# Hidden Markov modelling of spatio-temporal dynamics of measles in 1750–1850 Finland

### v1

@@ 3 Results
The model is estimated with MCMC using cmdstanr (Gabry and Češnovar,, 2022), which is an R interface (R Core Team,, 2023) for the probabilistic programming language Stan for statistical inference (Stan Development Team,, 2024). Posterior samples are drawn using NUTS sampler (Hoffman and Gelman,, 2014; Betancourt,, 2018) with four chains. Each chain consists of $7{,}500$ iterations, the first $5{,}000$ being discarded as warm-up. The computation with parallel chains takes approximately 30 hours. The estimation is done on a supercomputer node with four cores of Xeon Gold 6230 $2.1$ GHz processors, each core allocated $15$ GB of RAM. According to the MCMC diagnostics of the cmdstanr (Vehtari et al.,, 2021), the model converges without any divergences. The $\widehat{R}$ statistics are all below $1.004$, and the bulk effective sample sizes are roughly between $640$ and $28000$, the most inefficient one being $\varphi_{29,5}$, which corresponds to a town and month that, while not associated with a missing observation itself, is missing five out of six of the neighbouring observations. The R and Stan codes and the data used for the analysis are available on GitHub (https://github.com/tihepasa/spatialHMM).
@@
The spatial dimension of the states can be seen in Figure 3, which presents the estimated local probabilities and the corresponding data of observing at least one death caused by measles in each state. The probabilities are computed in accordance with the model, as in the temporal case above, aggregated over time, and displayed on a logarithmic scale to facilitate easier comparison both between and within the states. The smallest probabilities are associated with State 1, which seems to be common during the periods when the proportion of the observed deaths is low, see Figure 2. In States 1, 2 and 3, the probabilities seem to be larger in the southeastern parts of the country, and as moving from State 1 to the others the increased probabilities spread wider covering almost the whole study area in State 3. In State 4, the highest probabilities are located in the northern regions, and in State 5, they are concentrated in the southwestern half of the area.
@@
According to the transition probabilities, shown in Table 1, it is most likely to remain in the current state. If we disregard this possibility, the most probable transitions are from State 1 to State 2, from State 2 to State 1, from States 3 and 4 to State 2, and from State 5 to State 4. The estimates of the initial state probabilities $\rho_{s}$, along with other scalar parameters, can be found from Table 2. State 3 appears to be the starting point for the state trajectory in these data. However, the posterior distributions of the initial probabilities are quite wide. This is attributable to the fact that there is only one state trajectory to estimate, and an estimation based on a single observation does not provide much information.
@@
The local constants $\lambda_{i}$, independent of the state, are shown in Figure 4. As expected, there are no visible spatial patterns in these parameters aimed to capture spatially independent differences of the sites. The deviation $\sigma_{\lambda}$ of the constants $\lambda_{i}$ is 0.59, see also Table 2, suggesting that there is some non-spatial heterogeneity between the towns. For instance, in State 5, the 5% and 95% quantiles for the posterior mean of the local probabilities $\text{logit}^{-1}(\mu_{S}+\lambda_{i})$ are $(0.07,0.33)$ with median $0.17$.
@@
The spatial dependence in the spread of measles is accounted for by the spatial terms $\sigma_{\varphi}\varphi_{i,s}$. The deviation parameter $\sigma_{\varphi}$ is constant for all the sites and states and is estimated to be $0.89$, see also Table 2. The state-specific spatial terms $\sigma_{\varphi}\varphi_{i,s}$ are shown in Figure 5. The spatial patterns are similar to those of the estimated probabilities to observe at least one death caused by measles in Figure 3.
@@
To capture the yearly temporal variation, the model includes a seasonal component $\gamma_{t}$, which is shown in Figure 6. These monthly terms sum to zero over a year, i.e., they depict an average monthly variation additional to the other components of the model. The effect of the month increases from January to peak in May and decreases after that until December. The effect is positive, thereby increasing the probability of observing a death caused by measles, from March to July, and negative or decreasing the risk otherwise. A somewhat similar pattern, with a peak in May, was also identified in (Pasanen et al.,, 2023) for measles during 1820–1850 in Finland.
@@
From the sampled state trajectories of this left-to-right model, we computed the distribution of the most probable change point, which was identified as November 1812 (with a 95% posterior interval from September 1812 to February 1813). Table 3 displays the estimated emission matrix of this model. From this, we can conclude that the major changes were the increased probabilities of States 2, 3, and 5 and the decreased probability of State 1 after November 1812. This finding aligns well with the associated administrative changes that led to transferring the capital towards the east, and to altering the trade connections (Ojala and Räihä,, 2017).
@@
Determining the number of states in HMM is not a trivial task, and there is no definitive procedure for that, even though some suggestions exist (e.g., Pohle et al.,, 2017). In our case, we utilised five states, although a corresponding model was also estimated with four and six states. With six states there were several modes and convergence issues not only between but also within the chains. The model with four states, on the other hand, passed the convergence criteria but was not chosen here. Given that the states exhibited clear differences when there were five of them, we opted for the model with a greater number of hidden states. Presumably, this more detailed model also depicts the underlying phenomena more realistically in this context, since the epidemics are not discrete but continuous processes both in space and time.
For comparison, the supplementary material contains figures (Supplementary Figures 1–5) and tables (Supplementary Tables 1–2) for the four-state model, corresponding the ones represented here for the five-state model in Figures 2–6 and Tables 1–2. Based on Figures 2 and 3 and Supplementary Figures 1 and 2, it seems that States 1 and 2 of the five-state model have merged into State 1 of the four-state model, but otherwise the states correspond to each other.
@@ 4 Discussion
Using reported fatalities of measles in Finns throughout 1750–1850, and the Bayesian hidden Markov model with state-dependent spatial correlation structure, we identified five reoccurring epidemic states in Finland. The epidemics were characterised by two states of low burden of infection (States 1 and 2), and three states of higher infectious burden (States 3, 4 and 5), covering different parts of Finland (see Figure 3).
@@
Our analyses revealed also a distinctive change in the epidemic dynamics between November 1812 and February 1813, matching well with the changes caused by the war between Sweden and Russia over the area of Finland in 1808–1809, and the imposed large administrative changes, including the transfer of the capital from Turku to Helsinki, in 1812. These transformations strengthened the eastern influence and trade connections (Ojala and Räihä,, 2017). The most notable change in the estimated state trajectory after the annexation of Finland by Russia was the replacement of State 1 with a generally higher death observation probability of State 2 as a non-epidemic state. Also the likelihood of States 3 and 5 increased, reflecting the higher infection burden at southern Finland and Karelian Isthmus, or south-western Finland, respectively. The view of the strengthened eastern influence is further supported by the increase of the infection burden especially in the eastern towns during the low epidemic states (Figure 3), possibly indicating persistent infectious pressure from the closest city in the area, Saint Petersburg. Overall, infections became more common across whole Finland after the war.
@@
The dynamics of measles and other contagious infections are sensitive to population size (Keeling and Grenfell,, 1997). Due to the clear abrupt change in infection dynamics, we can perhaps safely rule out the role of gradual population increase (Voutilainen et al.,, 2020) here as the main driver of the change of the epidemics. Especially sparsely populated areas without endemic measles infections follow the changes of adjacent larger populations (Keeling and Grenfell,, 1997; Grenfell and Harwood,, 1997; Rohani et al.,, 2003; Ketola et al.,, 2021). Epidemics are affected by transport networks, geographical obstacles and borders (Vora et al.,, 2008). Finland and Sweden are separated by a large geographical barrier: The Gulf of Bothnia is 80–150 kilometers wide and 700 kilometers long, and obviously restricting the movement of people—and infections. Sea on the other side and the Russian border on the other side seem to have kept the infectious burden smaller in pre-war Finland compared with the post-war era when contacts to Russia and elsewhere increased (Ojala and Räihä,, 2017). During the history of nations, the borders have changed due to wars and the directions of transport and trade have shifted frequently, but the dynamics of epidemics due to such regime shifts have not been closely followed. Luckily, in our case, the administrative changes did not modify the data collection procedures of parishes so that we had data of the deaths available. This allowed us to study more closely the dynamics of epidemics due to changing transmission networks.
@@
When it comes to the local transmission routes inside Finland, the definition of the proximity between the towns plays a crucial role. Other relevant definitions for the neighbourhood could be road or water connections instead of or in addition to the border sharing we used. As during longer time periods the areal structure is rarely static due to wars and other environmental changes, it would be interesting to extend the model into a case in which the spatial structure and neighbours can vary in time. This would require some further specifications for the state-specific spatial coefficients, as allowing several spatial structures within one state could result in having not only $S$, but the number of different structures times $S$, states. To set the specifications and try this kind of approach in practice, one should have further information about the spatial structure and the changes in it, but technically it could be implemented with our model. Additionally, within a fixed spatial structure one could estimate if some neighbouring towns did not interact at all so that there were actually discontinuities, as in Balocchi and Jensen, (2019).
@@
As mentioned in section 1, there are several ways to model and estimate spatio-temporal connections. It would be natural to consider the phenomena as a continuous process instead of merely inspecting a collection of discrete states. One attractive way for this kind of more detailed analysis would be to fit a model without any hidden states, using random walks so that $y_{i,t}\sim\text{Bernoulli}(\mu_{t}+\lambda_{i}+\sigma_{\varphi}\varphi_{i,t}+\gamma_{t})$, where $\mu_{t}\sim\text{N}(\mu_{t-1},\sigma_{\mu}^{2})$, and $\boldsymbol{\varphi}_{t}\sim\text{N}(\boldsymbol{\varphi}_{t-1},(D-W)^{-1})$. In this case the spatio-temporal parameter $\varphi_{i,t}$ may be seen as an interaction parameter of the temporal and spatial terms, as described for example in Clayton and Bernardinelli, (1996). However, given the large number of time points, spatial units, and iterations needed for the convergence of the MCMC, this kind of model is computationally infeasible to estimate in our case. Additional downside would be that due to the parameter profusion, the results would not be as interpretable or offer such a straightforward insight as our hidden Markov model.
@@
We allowed for simplified, yet hopefully easy to understand, temporal dependency structure for the model via the parameters that depend on the state. Our method is especially convenient when there is a reason to suspect a presence of recurring patterns in the underlying phenomena. In other words, as we allow all potential transitions between our states so that it is possible to return to any of the previously occurred states later in time, also the latent process consists of repeated patterns. This enables finding the spatio-temporal epidemic dynamics and transmission routes illustrated by the repeated peak and fade-out patterns. In addition to the state-specific constants $\mu_{s}$ and spatial terms $\varphi_{i,t}$, also the local constant $\lambda_{i}$, deviation parameter $\sigma_{\varphi}$ and the seasonal term $\gamma_{t}$ could depend on state if seen appropriate. To avoid multimodality and convergence problems, we did not include those further complexities in our model.
@@
While our focus was on binary data, other observational distributions could be used and mixed, for example, a joint model of Poisson and Gaussian responses. There are previous indications of the profitability of the dichotomisation considering the last 30 years of our data (Pasanen et al.,, 2023). The model may also be extended to a case with exogenous covariates for the observational distributions or for the transition probabilities.

### v2

@@ 3 Results
The model is estimated with MCMC using cmdstanr (Gabry and Češnovar, 2022), which is an R interface (R Core Team, 2023) for the probabilistic programming language Stan for statistical inference (Stan Development Team, 2024). Posterior samples are drawn using NUTS sampler (Hoffman and Gelman, 2014; Betancourt, 2018) with four chains. Each chain consists of $7{,}500$ iterations, the first $5{,}000$ being discarded as warm-up. The computation with parallel chains takes approximately 30 hours. The estimation is done on a supercomputer node with four cores of Xeon Gold 6230 $2.1$ GHz processors, each core allocated $15$ GB of RAM. According to the MCMC diagnostics of the cmdstanr (Vehtari et al., 2021), the model converges without any divergences. The $\widehat{R}$ statistics are all below $1.004$, and the bulk effective sample sizes are roughly between $640$ and $28000$, the most inefficient one being $\varphi_{29,5}$, which corresponds to a town and month that, while not associated with a missing observation itself, is missing five out of six of the neighbouring observations. The R and Stan codes and the data used for the analysis are available on GitHub (https://github.com/tihepasa/spatialHMM).
@@
The spatial dimension of the states can be seen in Figure 3, which presents the estimated local probabilities and the corresponding data of observing at least one death caused by measles in each state. The probabilities are computed in accordance with the model, as in the temporal case above, aggregated over time, and displayed on a logarithmic scale to facilitate easier comparison both between and within the states. The smallest probabilities are associated with State 1, which seems to be common during the periods when the proportion of the towns observing at least one death is low, see Figure 2. In States 1, 2 and 3, the probabilities seem to be larger in the southeastern parts of the country, and as moving from State 1 to the others the increased probabilities spread wider covering almost the whole study area in State 3. In State 4, the highest probabilities are located in the northern regions, and in State 5, they are concentrated in the southwestern half of the area.
The dark blue areas in the data column of Figure 3 indicate that there are some towns where no deaths caused by measles were recorded, though it seems unlikely that any town would have been isolated enough to avoid the disease completely. Possible explanations for the lack of measles observations in these towns could be misdiagnoses or varying record-keeping practices (Pitkänen, 1977). Overall, when comparing the data and the predictions, there is approximately $0.3-3$ percentage point difference between the means of the data and the estimates. On average, the predictions underestimate the data, and the differences are smaller in the states where the probability to observe at least one death is smaller. This is not surprising as the discrete nature of the temporal dynamics of the model and the partial pooling of the site-specific variables smooth out some of the large observed mortality spikes.
@@
According to the transition probabilities, shown in Table 1, it is most likely to remain in the current state. If we disregard this possibility, the most probable transitions are from State 1 to State 2, from State 2 to State 1, from States 3 and 4 to State 2, and from State 5 to State 4. It should be remembered that these probabilities describe the overall transition probabilities instead of the likelinesses of the fast, abrupt monthly transitions. The estimates of the initial state probabilities $\rho_{s}$, along with other scalar parameters, can be found from Table 2. State 3 appears to be the starting point for the state trajectory in these data. However, the posterior distributions of the initial probabilities are quite wide. This is attributable to the fact that there is only one state trajectory to estimate, and an estimation based on a single observation does not provide much information.
@@
The local constants $\lambda_{i}$, independent of the state, are shown in Figure 4. As expected, there are no visible spatial patterns in these parameters aimed to capture spatially independent differences of the sites. The deviation $\sigma_{\lambda}$ of the constants $\lambda_{i}$ is 0.59 with posterior interval [0.53, 0.65], see also Table 2, suggesting that there is some non-spatial heterogeneity between the towns. For instance, in State 5, the 5% and 95% quantiles for the posterior mean of the local probabilities $\text{logit}^{-1}(\mu_{S}+\lambda_{i})$ are $(0.07,0.33)$ with median $0.17$.
@@
The spatial dependence in the spread of measles is accounted for by the spatial terms $\sigma_{\varphi}\varphi_{i,s}$. The deviation parameter $\sigma_{\varphi}$ is constant for all the sites and states and is estimated to be $0.89$ [0.82, 0.96], see also Table 2. The state-specific spatial terms $\sigma_{\varphi}\varphi_{i,s}$ are shown in Figure 5. The spatial patterns are similar to those of the estimated probabilities to observe at least one death caused by measles in Figure 3.
@@
To capture the yearly temporal variation, the model includes a seasonal component $\gamma_{t}$, which is shown in Figure 6. These monthly terms sum to zero over a year, i.e., they depict an average monthly variation additional to the other components of the model. The effect of the month increases from January to peak in May and decreases after that until December. The effect is positive, thereby increasing the probability of observing at least one death caused by measles, from March to July, and negative or decreasing the risk otherwise. A somewhat similar pattern, with a peak in May, was also identified in (Pasanen et al., 2024) for measles during 1820–1850 in Finland.
@@
From the sampled state trajectories of this left-to-right model, we computed the distribution of the most probable change point, which was identified as November 1812 (with a 95% posterior interval from September 1812 to February 1813). Table 3 displays the estimated emission matrix of this model. From this, we can conclude that the major changes were the increased probabilities of States 2, 3, and 5 and the decreased probability of State 1 after November 1812. This finding aligns well with the associated administrative changes that led to transferring the capital towards the east, and to altering the trade connections (Ojala and Räihä, 2017).
@@
Determining the number of states in HMM is not a trivial task, and there is no definitive procedure for that, even though some suggestions exist (e.g., Pohle et al., 2017). In our case, we utilised five states, although a corresponding model was also estimated with four and six states. With six states there were several modes and convergence issues not only between but also within the chains. The model with four states, on the other hand, passed the convergence criteria but was not chosen here. Given that the states exhibited clear differences when there were five of them, we opted for the model with a greater number of hidden states. Presumably, this more detailed model also depicts the underlying phenomena more realistically in this context, since the epidemics are not discrete but continuous processes both in space and time.
Additionally, we compared the four-state and five-state models using the expected log predictive density (ELPD), measuring the goodness of the entire predictive distribution (Vehtari et al., 2017). This was done leaving out one month and town at a time to estimate the ELPD with R package loo Vehtari et al. (2023). The results can be seen in Supplementary Table 3. Since a model with higher ELPD values corresponds to a model capable of greater posterior predictive accuracy for predicting new data points, and the ELPD of the five-state model is significantly higher in this case, we can say that the five-state model outperforms the four-state model, even though the ELPD is not completely reliable in our case based on a small proportion of high Pareto $\hat{k}$ values (Vehtari et al., 2017). The section Leave-one-out cross validation for hidden Markov models of the supplementary material contains also the derivation of the formula we use for computing the leave-one-out log-likelihoods used in LOO.
For comparison, the supplementary material contains figures (Supplementary Figures 1–5) and tables (Supplementary Tables 1–2) for the four-state model, corresponding the ones represented here for the five-state model in Figures 2–6 and Tables 1–2. Based on Figures 2 and 3 and Supplementary Figures 1 and 2, it seems that States 1 and 2 of the five-state model have merged into State 1 of the four-state model, but otherwise the states correspond to each other. Also in the four-state model a change in the dynamics towards the end of the time series is visible (see Supplementary Figure 1).
@@ 4 Discussion
Using reported fatalities of measles in Finland throughout 1750–1850, and the Bayesian hidden Markov model with state-dependent spatial correlation structure, we identified five reoccurring epidemic states in Finland. The epidemics were characterised by two states of low burden of infection (States 1 and 2), and three states of higher infectious burden (States 3, 4 and 5), covering different parts of Finland (see Figure 3).
@@
Our analyses revealed also a distinctive change in the epidemic dynamics between November 1812 and February 1813, matching well with the changes caused by the war between Sweden and Russia over the area of Finland in 1808–1809, and the imposed large administrative changes, including the transfer of the capital from Turku to Helsinki, in 1812. These transformations strengthened the eastern influence and trade connections (Ojala and Räihä, 2017). The most notable change in the estimated state trajectory after the annexation of Finland by Russia was the replacement of State 1 with a generally higher death observation probability of State 2 as a non-epidemic state. Also the likelihood of States 3 and 5 increased, reflecting the higher infection burden at southern Finland and Karelian Isthmus, or south-western Finland, respectively. The view of the strengthened eastern influence is further supported by the increase of the infection burden especially in the eastern towns during the low epidemic states (Figure 3), possibly indicating persistent infectious pressure from the closest city in the area, Saint Petersburg. Overall, infections became more common across whole Finland after the war.
@@
The dynamics of measles and other contagious infections are sensitive to population size (Keeling and Grenfell, 1997). Due to the clear abrupt change in infection dynamics, we can perhaps safely rule out the role of the gradual population increase (Voutilainen et al., 2020) here as the main driver of the change of the epidemics. Further examination considering particularly the spatial effect of the population sizes remains as a future work due to the missing local population size information. Especially sparsely populated areas without endemic measles infections follow the changes of adjacent larger populations (Keeling and Grenfell, 1997; Grenfell and Harwood, 1997; Rohani et al., 2003; Ketola et al., 2021). Epidemics are affected by transport networks, geographical obstacles and borders (Vora et al., 2008). Finland and Sweden are separated by a large geographical barrier: The Gulf of Bothnia is 80–150 kilometers wide and 700 kilometers long, and obviously restricting the movement of people—and infections. Sea on the other side and the Russian border on the other side seem to have kept the infectious burden smaller in pre-war Finland compared with the post-war era when contacts to Russia and elsewhere increased (Ojala and Räihä, 2017). During the history of nations, the borders have changed due to wars and the directions of transport and trade have shifted frequently, but the dynamics of epidemics due to such regime shifts have not been closely followed. Luckily, in our case, the administrative changes did not modify the data collection procedures of parishes so that we had data of the deaths available. This allowed us to study more closely the dynamics of epidemics due to changing transmission networks.
@@
When it comes to the local transmission routes inside Finland, the definition of the proximity between the towns plays a crucial role. Other relevant definitions for the neighbourhood could be road or water connections instead of or in addition to the border sharing we used. As during longer time periods the areal structure is rarely static due to wars and other environmental changes, it would be interesting to extend the model into a case in which the spatial structure and neighbours can vary in time. This would require some further specifications for the state-specific spatial coefficients, as allowing several spatial structures within one state could result in having not only $S$, but the number of different structures times $S$, states. To set the specifications and try this kind of approach in practice, one should have further information about the spatial structure and the changes in it, but technically it could be implemented with our model. Additionally, within a fixed spatial structure one could estimate if some neighbouring towns did not interact at all so that there were actually discontinuities, as in Balocchi and Jensen (2019).
@@
As mentioned in Section 1, there are several ways to model and estimate spatio-temporal connections. It would be natural to consider the phenomena as a continuous process instead of merely inspecting a collection of discrete states. One attractive way for this kind of more detailed analysis would be to fit a model without any hidden states, using random walks so that $y_{i,t}\sim\text{Bernoulli}(\mu_{t}+\lambda_{i}+\sigma_{\varphi}\varphi_{i,t}+\gamma_{t})$, where $\mu_{t}\sim\text{N}(\mu_{t-1},\sigma_{\mu}^{2})$, and $\boldsymbol{\varphi}_{t}\sim\text{N}(\boldsymbol{\varphi}_{t-1},(D-W)^{-1})$. In this case the spatio-temporal parameter $\varphi_{i,t}$ may be seen as an interaction parameter of the temporal and spatial terms, as described for example in Clayton and Bernardinelli (1996). Given the large number of time points, spatial units, and iterations needed for the convergence of the MCMC, this kind of model is computationally infeasible to estimate in our case using MCMC, although approximate solutions using INLA would have likely been obtainable. However, the downside of such models is that due to the parameter profusion, the results would not be as interpretable or offer such a straightforward insight as our hidden Markov model.
@@
We allowed for simplified, yet hopefully easy to understand, temporal dependency structure for the model via the parameters that depend on the state. Our method is especially convenient when there is a reason to suspect a presence of recurring patterns in the underlying phenomena. In other words, as we allow all potential transitions between our states so that it is possible to return to any of the previously occurred states later in time, also the latent process consists of repeated patterns. This enables finding the spatio-temporal epidemic dynamics and transmission routes illustrated by the repeated peak and fade-out patterns. In addition to the state-specific constants $\mu_{s}$ and spatial terms $\varphi_{i,t}$, also the local constant $\lambda_{i}$, deviation parameter $\sigma_{\varphi}$ and the seasonal term $\gamma_{t}$ could depend on state if seen appropriate. Our attempts to include the state dependent local constants $\lambda_{i}$ and deviations $\sigma_{\varphi}$ led to multimodality and convergence issues. The same was also true for an alternative model in which, instead of the state-dependent $\mu_{s}$, the model included a temporal trend $\mu_{t}$ which evolved continuously over time (in which case the states describe only changes in the spatial patterns). Thus, the inclusion of these further complexities in our model remains as a future work.
@@
While our focus was on binary data, other observational distributions could be used and mixed, for example, a joint model of Poisson and Gaussian responses. There are previous indications of the profitability of the dichotomisation considering the last 30 years of our data (Pasanen et al., 2024). The model may also be extended to a case with exogenous covariates for the observational distributions or for the transition probabilities.